OpenStudy (anonymous):

Find the constant sum for an ellipse with foci F1 (2, 0), F2 (-6, 0) and the point on the ellipse (2, 6). The constant sum is: ____. Enter only a number.

OpenStudy (nurali):

One property of an ellipse is that, from any point, the sum of the distances to the two foci is constant. √((2- 2)² + (6 - 0)²) + √((2 - (-6))² + (6 - 0)²) = √36 + √100 = 6 + 10 = 16

OpenStudy (anonymous):

oh ok yes i understand thanks :)